Copper contains 8.4 Times 10^28 free electrons per
Answer #1

given are

N=8.4*10^28 electrons

A=7.4*10^-7 m^2

I=1amp

using formula

I = nAve

v = I/(nAe)

v=1/8.4*10^28*7.4*10^-7*1.6*10^-19

v=0.01005*10^-2m/sec